OpenStudy (anonymous):

lim x->4 = (1/sqrt(x)-2) - (4/x-4)

OpenStudy (anonymous):

\[\text{Limit}\left[\left(\frac{1}{\sqrt{x}}-2\right)-\left(\frac{4}{x}-4\right),x\to 4\right]=\frac{3}{2} \]